# Long Chain-of-Thought for Reasoning Large Language Models

### v4

@@ 1 Introduction
In recent years, as shown in Figure 1, the emergence of reasoning large language models (RLLMs) such as OpenAI o1 [289] and DeepSeek R1 [213] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [668, 943, 694, 75, 925, 187, 1071], even passing Turing Test [315]. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [1104, 593, 596, 573]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by inference-time scaling [398, 706, 501]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [597, 526].
@@ 2.2.2 Extensive Exploration for Long CoT
As shown by Figure 2, Long CoT encourages branching out to extensively explore uncertain or unknown logical nodes, thereby expanding the potential set of reasoning paths. This exploration is particularly critical when solving problems characterized by ambiguity, incomplete information, or multiple possible solutions. More specifically, we describe how extensive exploration primarily addresses the relaxation of the second constraint in Equation 1, which can be formalized as follows:
@@ 2.2.3 Feasible Reflection for Long CoT
$\widetilde{n}_{i+1}\leftarrow\text{Refine}(n_{i+1}|\texttt{CoT}_{L}^{i},\mathcal{F}_{i},n_{j})$ (6)
@@ 3.1.1 Long CoT External Behavior Analysis
Recent advances in inference-time scaling algorithms [501, 811] have garnered significant attention, particularly for their ability to extend reasoning length and improve performance [501, 432, 841]. Specifically, Brown et al. [53] identify a phenomenon called "Large Language Monkeys," in which a series of reasoning tasks show that with enough trials, a correct result can be achieved. Additionally, o1 [289] and R1 [213] demonstrated that directly scaling the length of model inference improves final performance.
@@
As RLLMs evolve, it is crucial to navigate the debate around the selection between process supervision and outcome supervision, two key reinforcement learning paradigms for complex reasoning tasks. The phenomenon of choosing between these two approaches has become a pivotal issue, as it is essential to differentiate and decide which supervision strategy is more suitable for specific tasks [865, 175, 1019]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two approaches remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [1060, 652]. Additionally, PRM faces the issue of reward hacking [12, 144, 547, 29, 379], where agents exploit flaws in the reward function to produce unintended behaviors [213]. Addressing this to surpass rule-based reward systems has become an important research area [213, 852, 568]. Furthermore, Lampinen et al. [349] and Tan [680] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [298] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors.
@@ 3.1.2 Long CoT Internal Mechanism Analysis
Recent studies have explored the internal mechanisms underlying the coherent rationale outputs of Long CoT, with particular emphasis on attention mechanisms [648, 606]. These studies primarily examine neural substructures in RLLMs, framing CoT reasoning from a white-box perspective [789, 954, 219, 159]. Weston and Sukhbaatar [817] introduces the concept of System 2 Attention (S2A), which demonstrates Long CoT generation by selectively focusing attention on relevant information. Additionally, Li et al. [387] explore gradient distributions between direct output and Long CoT layers, revealing that Long CoT layers help maintain stability by distinguishing relevant from irrelevant reasoning. Finally, Zhang et al. [1028] conceptualize RLLMs as finite state automata, offering further insight into how internal dynamics influence external behavior. Despite Short CoT’s struggles with self-correction, Bertolazzi et al. [43] show that these models rely on consistency heads (attention heads) to assess the alignment of numerical values in arithmetic solutions through internal shortcuts.
@@ 3.2.3 Benchmarks
  • Scientific Reasoning: Scientific Reasoning benchmarks, such as GPQA Diamond [611], MMLU-Pro [791], and SuperGPQA [155], assess multi-domain reasoning in fields like chemistry, biology, and physics. These benchmarks test models’ ability to not only accumulate knowledge but also integrate it for problem-solving. Humanity’s Last Exam (HLE) [576] further challenges models by requiring deep interdisciplinary reasoning across scientific disciplines. Further, Chung et al. [134] propose TPBench to evaluate the effectiveness of RLLMs in solving theoretical physics problems.
  • Medical Reasoning: In the realm of Medical Reasoning, the need for complex, domain-specific, and accurate reasoning is paramount [1052, 986, 871, 563]. Benchmarks, such as MedQA[309], JAMA Clinical Challenge [72], and Medbullets [72], simulate diagnostic and treatment decision-making processes, reflecting real-world medical practice. These benchmarks evaluate a model’s handling of medical knowledge and reasoning, from diagnosis to treatment planning. Additionally, MedXpertQA [1107] introduces a comprehensive evaluation framework combining text and multimodal data, specifically assessing AI’s reasoning capabilities in healthcare.
@@ 3.2.5 Advanced Evaluation
Agentic and Embodied reasoning requires models to demonstrate an understanding of real-world interactions, tool use, and adaptive reasoning in response to change. To assess real-world understanding, Wang et al. [768] introduce a benchmark that evaluates agents’ ability to reason about physical concepts. Zhang et al. [1024] extend this by assessing agents’ interactions with real-world physics. Additionally, realistic tasks often demand complex planning and tool usage, necessitating benchmarks to evaluate agent reasoning. These benchmarks assess agents’ abilities to navigate and complete tasks in digital environments. Building on this, Huang et al. [266] propose a framework for evaluating decision-making in multi-agent, competitive settings. Nath et al. [536] introduce ToolComp, a benchmark designed to evaluate multi-step tool-use reasoning. To analyze adaptive reasoning in the face of real-world change, OSWorld [853], CogAgent [243], Mobile-Agent-E [798], WebShop [918], WebArena [1084], and WebGames [708] assess AI systems across domains such as operating systems, mobile GUIs, browser tasks, and interactive entertainment [1064, 752, 489]. Hu et al. [255] present Text2World, which evaluates agents’ ability to generate interactive environments from text to test agent adaptability [957].
@@
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [484], M3CoT-Science [86], and ScienceBoard [670] evaluate how well models combine science information with Long CoT reasoning across various scientific domains [930]. Further, Guo et al. [215] propose MolPuzzle for the evaluation of molecular structure elucidation.
  • Commonsense Puzzle: This area focuses on commonsense reasoning, where systems combine reasoning cues and images to make deeper conclusions. Chen et al. [86] introduce M3CoT-Commensense, which incorporates commonsense Long CoT reasoning for complex multimodal interactions. Further, PuzzleVQA [123] and LEGO-Puzzles [683] focus more on abstract and spatial puzzle reasoning, respectively. Additionally, Wang et al. [733] propose two benchmarks: Clue-Visual Question Answering (CVQA), which tests visual comprehension through three task types, and Clue of Password-Visual Question Answering (CPVQA), which features two task types focusing on the interpretation and application of visual data.
@@ 4.1.1 Natural Language Deep Reasoning
Traditionally, researchers have sought to adapt natural language for intuitive and free-flowing deep reasoning [806, 1076, 285, 591, 1030, 738, 192]. Early work by Wei et al. [806] demonstrated that the use of natural language Long CoT significantly enhances the reasoning capabilities of RLLMs. Further, the Natural Program framework [437] allows RLLMs to engage in deeper natural language reasoning by ensuring a more structured and rigorous logical analysis. More recently, CodeI/O [381] has introduced a technique that reorganizes code-based reasoning patterns into natural language formats, further boosting the reasoning potential of RLLMs.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [657, 143, 614, 305]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [780, 975] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces [914, 970]. Further, Coconut [222] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [1029, 678]. At the extreme, Heima [635] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [337] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of inference-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Geiping et al. [191] and Saunshi et al. [620] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block” [488]. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [104] leverages the original transformer layer as a recurrent “thought block”, selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens. Further, System-1.5 Reasoning [778] defines two dynamic shortcuts. The Model Depth Shortcut (DS) lets non-critical tokens exit early via lightweight adapter branches while routing critical tokens through deeper Transformer layers, thus supporting adaptive, vertical reasoning. The Step Shortcut (SS) reuses hidden states across decoding steps to bypass trivial iterations and enable horizontal reasoning in latent space.
@@ 5 Feasible Reflection for Long CoT
Feasible Reflection is a pivotal component of Long CoT reasoning, enabling LLMs to handle complex tasks through iterative feedback and refinement [386, 179]. Specifically, it comprises two primary stages: (1) Feedback (§ 5.1), which generates feedback signals $\mathcal{F}_{i}$ to correct node $n_{j}$ in Equation (5); and (2) Refinement (§ 5.2), which adjusts the subsequent node $n_{i+1}$ according to the feedback in Equation (LABEL:eq:refinement).
@@ 5.1.1 Overall Feedback
Research on feedback from RLLMs centers on detecting errors and biases through natural language feedback, also known as LLM-as-Judge, self-reflection or self-critique [257, 317, 28, 612, 525, 772, 964, 861]. This method has led to significant improvements across various tasks, particularly in self-correction [816, 1067, 193, 172, 1034]. Huang et al. [269] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [619, 378]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [929, 36, 37, 265].
@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [128, 402, 505]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [426] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [763] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [69, 963]. Following this, methods like QwQ [703], Skywork-o1 [544], AceMath [477], and PRIME [137] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [998] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [869] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [977] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [205] and Zhang et al. [1033] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. Further, Wang et al. [777] propose VisualPRM400K and expand this paradigm to multimodal scenarios. (2) Outcome Annotated PRM Training: Alternative approaches, such as ReST-MCTS* [994], OVM [942], Implicit PRM [962], AutoPSV [483], and DVO [1000], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [857, 617]. UAS [944] incorporates uncertainty-aware value models [258] into feedback predictions [472, 157, 910]. Additionally, Aurora [682] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [624] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [897], Lee et al. [357], Yoon et al. [938] extend these paradigms to the token level. Moreover, Chen et al. [105] expand these into interactive agent scenarios, allowing for automatically learning reward models from the environment without additional manual annotation. Wang et al. [802] equip a dual-layer MLP module to evaluate the reward at each step, successfully integrating the policy model and PRM into a unified interface without additional process annotations, reducing over 99% of PRM parameters for efficient reasoning.
@@
Dyve-14B [1069] DeepSeek-R1-Distill-Qwen-14B [213] 68.5 58.3 49.0 47.2 - - -

### v5

@@ 1 Introduction
In recent years, as shown in Figure 1, the emergence of reasoning large language models (RLLMs) such as OpenAI o1 [307] and DeepSeek R1 [227] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [696, 980, 722, 79, 961, 200, 1113, 793], even passing Turing Test [334]. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [1147, 619, 622, 599]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by inference-time scaling [419, 733, 524]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [623, 550].
@@ 2.2.2 Extensive Exploration for Long CoT
As shown by Figure 2, Long CoT encourages branching out to extensively explore uncertain or unknown logical nodes, thereby expanding the potential set of reasoning paths. This exploration is particularly critical when solving problems characterized by ambiguity, incomplete information, or multiple possible solutions [43, 1016, 871]. More specifically, we describe how extensive exploration primarily addresses the relaxation of the second constraint in Equation 1, which can be formalized as follows:
@@ 2.2.3 Feasible Reflection for Long CoT
$\widetilde{n}_{i+1}\leftarrow\text{Refine}(n_{i+1}|\texttt{CoT}_{L}^{i},\mathcal{F}_{i},n_{j}),$ (6)
@@ 3.1.1 Long CoT External Behavior Analysis
Recent advances in inference-time scaling algorithms [524, 843] have garnered significant attention, particularly for their ability to extend reasoning length and improve performance [524, 455, 875]. Specifically, Brown et al. [57] identify a phenomenon called “Large Language Monkeys”, in which a series of reasoning tasks show that with enough trials, a correct result can be achieved. Additionally, o1 [307] and R1 [227] demonstrated that directly scaling the length of model inference improves final performance.
@@
As RLLMs evolve, it is crucial to navigate the debate around the selection between process supervision and outcome supervision, two key reinforcement learning paradigms for complex reasoning tasks. The phenomenon of choosing between these two approaches has become a pivotal issue, as it is essential to differentiate and decide which supervision strategy is more suitable for specific tasks [899, 187, 1059]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two approaches remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [1102, 679]. Additionally, PRM faces the issue of reward hacking [13, 152, 573, 30, 399], where agents exploit flaws in the reward function to produce unintended behaviors [227]. Addressing this to surpass rule-based reward systems has become an important research area [227, 886, 594]. Furthermore, Lampinen et al. [368] and Tan [708] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [317] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors. More strictly, He et al. [247] mathematically demonstrate that outcome-level rewards suffice for online reinforcement learning in RLLMs.
@@
Reinforcement Learning Entropy Phenomenon
In reinforcement learning for Long CoT, the entropy mechanism is a crucial factor influencing the performance of RLLMs. Policy entropy measures the diversity and exploratory strength of a model’s outputs. By managing this entropy effectively, a model preserves exploration and thus excels on complex reasoning tasks. Earlier, Jang and Kim [310] investigate how initial entropy affects exploration in deep RL and proposed an entropy-aware initialization to encourage effective exploration. Building on this, Zhang et al. [1036] developed an Entropy-Regularized PRM that balances policy updates against large deviations from the starting distribution, thereby improving reasoning. Cheng et al. [116] found that high-entropy regions correlate positively with three exploratory reasoning behaviors: (1) key tokens linking logical steps, (2) self-verification and correction, and (3) rare behaviors underrepresented in the base model. Most recently, Agarwal et al. [5] introduced an Entropy Minimization method and demonstrated its strong impact on LLM performance in mathematical, physical, and coding tasks.
However, recent research indicates that, during early training, policy entropy declines sharply, causing the model to converge prematurely on specific output patterns and limiting further reasoning improvement [144]. In reinforcement learning, policy entropy ($H$) and downstream task performance ($R$) follow an exponential relation: $R=-a\cdot e^{H}+b,$ so a drop in entropy produces a rapid performance decline until saturation. This “policy entropy collapse” is common without entropy control, as reduced entropy constrains exploration and stalls reasoning gains [144]. To counter this collapse, two methods, Clip-Cov and KL-Cov, regulate entropy by constraining updates on high-covariance tokens. Clip-Cov clips their update magnitudes, whereas KL-Cov imposes a Kullback–Leibler penalty. Empirical results show both techniques prevent collapse and enhance reasoning performance [144].
@@ 3.1.2 Long CoT Internal Mechanism Analysis
Recent studies have explored the internal mechanisms underlying the coherent rationale outputs of Long CoT, with particular emphasis on attention mechanisms [675, 632]. These studies primarily examine neural substructures in RLLMs, framing CoT reasoning from a white-box perspective [819, 992, 233, 169]. Weston and Sukhbaatar [849] introduces the concept of System 2 Attention (S2A), which demonstrates Long CoT generation by selectively focusing attention on relevant information. Additionally, Li et al. [407] explore gradient distributions between direct output and Long CoT layers, revealing that Long CoT layers help maintain stability by distinguishing relevant from irrelevant reasoning [840]. Finally, Zhang et al. [1068] conceptualize RLLMs as finite state automata, offering further insight into how internal dynamics influence external behavior. Despite Short CoT’s struggles with self-correction, Bertolazzi et al. [47] show that these models rely on consistency heads (attention heads) to assess the alignment of numerical values in arithmetic solutions through internal shortcuts.
@@ 3.2.3 Benchmarks
  • Scientific Reasoning: Scientific Reasoning benchmarks, such as GPQA Diamond [637], MMLU-Pro [821], and SuperGPQA [165], assess multi-domain reasoning in fields like chemistry, biology, and physics [157]. These benchmarks test models’ ability to not only accumulate knowledge but also integrate it for problem-solving. Humanity’s Last Exam (HLE) [602] further challenges models by requiring deep interdisciplinary reasoning across scientific disciplines. Further, Chung et al. [140] propose TPBench to evaluate the effectiveness of RLLMs in solving theoretical physics problems.
  • Medical Reasoning: In the realm of Medical Reasoning, the need for complex, domain-specific, and accurate reasoning is paramount [1094, 1024, 905, 589]. Benchmarks, such as MedQA [328], JAMA Clinical Challenge [76], LLMEval-Med [1049] and Medbullets [76], simulate diagnostic and treatment decision-making processes, reflecting real-world medical practice. These benchmarks evaluate a model’s handling of medical knowledge and reasoning, from diagnosis to treatment planning. Additionally, MedXpertQA [1150] introduces a comprehensive evaluation framework combining text and multimodal data, specifically assessing AI’s reasoning capabilities in healthcare.
@@ 3.2.5 Advanced Evaluation
Agentic and Embodied reasoning requires models to demonstrate an understanding of real-world interactions, tool use, and adaptive reasoning in response to change. To assess real-world understanding, Wang et al. [798] introduce a benchmark that evaluates agents’ ability to reason about physical concepts. Zhang et al. [1064] extend this by assessing agents’ interactions with real-world physics. Additionally, realistic tasks often demand complex planning and tool usage, necessitating benchmarks to evaluate agent reasoning. These benchmarks assess agents’ abilities to navigate and complete tasks in digital environments. Building on this, Huang et al. [283] propose a framework for evaluating decision-making in multi-agent, competitive settings. Nath et al. [562] introduce ToolComp, a benchmark designed to evaluate multi-step tool-use reasoning. To analyze adaptive reasoning in the face of real-world change, OSWorld [887], CogAgent [260], Mobile-Agent-E [828], WebShop [954], WebArena [1126], WGSR-Bench [972], and WebGames [735] assess AI systems across domains such as operating systems, mobile GUIs, browser tasks, and interactive entertainment [1106, 780, 512, 552]. Hu et al. [272] present Text2World, which evaluates agents’ ability to generate interactive environments from text to test agent adaptability [995].
@@
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [507], M3CoT-Science [91], BMMR [874], and ScienceBoard [698] evaluate how well models combine science information with Long CoT reasoning across various scientific domains [966]. Further, Guo et al. [229] propose MolPuzzle for the evaluation of molecular structure elucidation.
  • Commonsense Puzzle: This area focuses on commonsense reasoning, where systems combine reasoning cues and images to make deeper conclusions. Chen et al. [91] introduce M3CoT-Commensense, which incorporates commonsense Long CoT reasoning for complex multimodal interactions. Further, PuzzleVQA [128], MMReason [953] and LEGO-Puzzles [711] focus more on abstract and spatial puzzle reasoning, respectively. Additionally, Wang et al. [760] propose two benchmarks: Clue-Visual Question Answering (CVQA), which tests visual comprehension through three task types, and Clue of Password-Visual Question Answering (CPVQA), which features two task types focusing on the interpretation and application of visual data.
@@ 4.1.1 Natural Language Deep Reasoning
Traditionally, researchers have sought to adapt natural language for intuitive and free-flowing deep reasoning [836, 1118, 303, 617, 1070, 765, 205]. Early work by Wei et al. [836] demonstrated that the use of natural language Long CoT significantly enhances the reasoning capabilities of RLLMs. Further, the Natural Program framework [460] allows RLLMs to engage in deeper natural language reasoning by ensuring a more structured and rigorous logical analysis. More recently, CodeI/O [401] has introduced a technique that reorganizes code-based reasoning patterns into natural language formats, further boosting the reasoning potential of RLLMs [36]. Similarly, Li et al. [387] propose CoRT, which integrates code into reasoning to facilitate a mixture of formats, resulting in improved cognitive performance.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [684, 151, 640, 324]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [810, 1013] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces [949, 1008]. Further, Coconut [236] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [1069, 706]. At the extreme, Heima [662] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [356] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of inference-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Schöne et al. [647], Geiping et al. [204], and Saunshi et al. [646] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block” [511]. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [109] leverages the original transformer layer as a recurrent “thought block”, selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens. Further, System-1.5 Reasoning [808] defines two dynamic shortcuts. The Model Depth Shortcut (DS) lets non-critical tokens exit early via lightweight adapter branches while routing critical tokens through deeper Transformer layers, thus supporting adaptive, vertical reasoning. The Step Shortcut (SS) reuses hidden states across decoding steps to bypass trivial iterations and enable horizontal reasoning in latent space.
@@ 5 Feasible Reflection for Long CoT
Feasible Reflection is a pivotal component of Long CoT reasoning, enabling LLMs to handle complex tasks through iterative feedback and refinement [406, 192]. Specifically, it comprises two primary stages: (1) Feedback (§ 5.1), which generates feedback signals $\mathcal{F}_{i}$ to correct node $n_{j}$ in Equation (5); and (2) Refinement (§ 5.2), which adjusts the subsequent node $n_{i+1}$ according to the feedback in Equation (6).
@@ 5.1.1 Overall Feedback
Research on feedback from RLLMs centers on detecting errors and biases through natural language feedback, also known as LLM-as-Judge, self-reflection or self-critique [274, 336, 29, 638, 549, 802, 1002, 895, 529]. This method has led to significant improvements across various tasks, particularly in self-correction [848, 1109, 206, 184, 1075]. Huang et al. [286] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [645, 398]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [965, 39, 40, 282].
@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [134, 423, 528]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [449] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [792] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [73, 1001]. Following this, methods like QwQ [731], Skywork-o1 [570], AceMath [500], and PRIME [143] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [1036] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [903] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [1015] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [219] and Zhang et al. [1074] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. Further, Wang et al. [807] propose VisualPRM400K and expand this paradigm to multimodal scenarios. (2) Outcome Annotated PRM Training: Alternative approaches, such as ReST-MCTS* [1032], OVM [979], Implicit PRM [1000], AutoPSV [506], and DVO [1038], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [891, 643]. UAS [981] incorporates uncertainty-aware value models [275] into feedback predictions [495, 167, 945, 1089]. Additionally, Aurora [710] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [651] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [932], Lee et al. [376], Yoon et al. [975] extend these paradigms to the token level. Moreover, Chen et al. [110] expand these into interactive agent scenarios, allowing for automatically learning reward models from the environment without additional manual annotation. Wang et al. [832] equip a dual-layer MLP module to evaluate the reward at each step, successfully integrating the policy model and PRM into a unified interface without additional process annotations, reducing over 99% of PRM parameters for efficient reasoning.
@@
Dyve-14B [$1$] DeepSeek-R1-Distill-Qwen-14B [227] 68.5 58.3 49.0 47.2 - - -